A company currently pays a dividend of $4 per share (D0= $4). It is estimated that the company’s dividend will grow at a rate of 10% per year for the next 2 years, and then at a constant rate of 5% thereafter. The company’s stock has a beta of 6, the risk-free rate is 4% and the market risk premium is 2%. What is your estimate of the stock’s current price? Please solve in Excel. Thank you!
Answer #1

First we calculate the required return as per CAPM:

return=risk free+ (Beta*Market risk premium)
Market risk premium=Market return-risk free rate)

Return=4+(6*2)=16%

Dividend year 1=4.00*1.10=4.40

Year Div Terminal value Total Discount factor Present Value
1 $ 4.4000 $     4.40 0.847457627 $                 3.73
2 $ 4.8400 $            46.2000 $   51.04 0.71818443 $               36.66
3 $ 5.0820 0.608630873 $                      -  
4 Total $          40.3849

Discount factors (DF)are calculated as:1/(1+r)^n

Where r is required return and n is the year

For the perpetual growth period, terminal value (TV) of dividends = Dividend next period/(r-g) where r is the required return and g is the growth rate.

Terminal value of dividends= 5.082/(0.16-0.05)=$46.20